User name C. Fernandez-Ramirez
Log entry time 19:10:28 on December 9,2007
Entry number 219412
keyword=Return to beam permit
Left Dipole and Q1 were reset from downstairs. Scot had to go down twice. He also fixed the meter of Helium because it was giving a false feedback.